A hydrocarbon was found to contain 0.96g of carbon and 0.2g of hydrogen. Calculate the simplest formula


1
Expert's answer
2021-11-29T04:40:10-0500

Moles of Carbon = "\frac{0.95}{12}=0.08moles"


Moles of Hydrogen "=\frac{0.2}{1}=0.2moles"


Mole ratios "Carbon= \frac{0.08}{0.08}=1" "Hydrogen=\frac{0.2}{0.08}=2.5"


"CH_3"